# QE4PE: Word-level Quality Estimation for Human Post-Editing

### v1

@@ 3.4 Participants
For both directions, professional translation providers Translated and Global Textware recruited three translators for the Oracle post-editing stage, 12 for Pre, Main and Post tasks, and finally six for the QA stage. All translators were freelancers with native proficiency in their target language and self-assessed proficiency of at least C1 in English. Almost all translators had more than two years of professional translation experience and regularly post-edited MT outputs (details in Table 7).
@@ 3.5 Editing Interface
We develop a custom interface, which we name Groningen Translation Environment (GroTE, Figure 2), to support editing over texts with word-level highlights. While the MMPE tool used by Shenoy et al. (2021) provide extensive multimodal functionalities (Herbig et al., 2020), we aim for a bare-bones setup to avoid confounders in the evaluation. GroTE is a web interface based on Gradio (Abid et al., 2019) and hosted on the Hugging Face Spaces to enable multi-user data collection online. Upon loading a document, source texts and MT outputs for all segments are presented in two columns following standard industry practices. For modalities with highlights, the interface provides an informative message and supports the removal of all highlights on a segment via a button, with highlights on words disappearing automatically upon editing, as in Shenoy et al. (2021). The interface supports real-time logging of user actions to analyze the editing process. In particular, we log the start and end times for each edited document, the accessing and exiting of segment textboxes, highlight removals, and individual keystrokes during editing.
GroTE intentionally lacks standard features such as translation memories, glossaries, and spellchecking to ensure equal familiarity among translators, ultimately controlling for editor proficiency with these tools, as done in previous studies (Shenoy et al., 2021; Sarti et al., 2022). While most translators noted the lack of features in our usability assessment, the majority also found the interface easy to set up, access, and use (Table 7).
@@ 4.1 Productivity
We obtain segment- and document-level edit times and compute editing productivity as the number of processed source characters per second over the sum of all document-level edit times. To account for potential breaks taken by post-editors during editing, we filter out pauses between logged actions longer than 5 minutes. We note that our findings are not sensitive to this filtering.
@@
Figure 3 shows translators’ productivity across stages, with every dot corresponding to the productivity of a single individual. We observe that no highlight modality leads to systematically faster editing across all speed groups and that the ordering of Pre-task speed groups is maintained in the following stages despite the different highlight modalities. These results suggest that individual variability in editing speed is more critical than highlight modality in predicting editing speed. However, faster EnglishDutch translators achieve outstanding productivity ( char/min) only for No Highlight, and, Oracle modalities, suggesting that lower-quality highlights hinder editing speed.
We validate these observations by fitting a negative binomial mixed-effect model on segment-level editing times (model details in Table 5). Excluding random factors such as translator and segment identity from the model produces a significant drop in explained variance, confirming the inherent variability of editing times (). Model coefficients show that MT output length and the proportion of highlighted characters are the main factors driving an increase in editing times, possibly reflecting an increase in cognitive effort to process additional information. We find highlights to have a significant impact on increasing the editing speed of EnglishItalian translators , but a minimal impact for EnglishDutch. Comparing the productivity of the same translator editing with and without highlights (Main vs Post), two-thirds of the translators editing with highlights were up to two times slower on biomedical texts. However, the same proportion of translators was up to three times faster on social media texts across both directions.
@@ 4.2 Highlights and Edits
First, we quantify how different modalities agree in terms of highlights’ distribution and editing. We find that cross-modality highlight overlap ranges between 15% and 39%, with the highest overlap for EnglishItalian social media and EnglishDutch biomedical texts.444Scores are normalized to account for highlight frequencies across modalities. Highlight agreement shown in Table 8. Despite the relatively low highlight agreement, we find an average agreement of 73% for post-edited characters across modalities. This suggests that edits are generally uniform regardless of highlight modalities and are not necessarily restricted to highlighted spans.555Editing agreement is shown in Figure 8.
@@
Table 1 (Base Freq.) shows raw highlight and edit frequencies across modalities. We observe different trends across the two language pairs: for EnglishItalian, post-editors working with highlights edit more than twice as much as translators with No Highlight, regardless of the highlight modality. On the contrary, for EnglishDutch they edit 33% less in the same setting. These results suggest a different attitude towards acting upon highlighted potential issues across the two translation directions, with EnglishItalian translators appearing to be conditioned to edit more when highlights are present. We introduce four metrics to quantify highlights-edits overlap:
@@
Intuitively, character-level recall should be more indicative of highlight quality compared to precision , provided that word-level highlights can be useful even when not minimal.666For example, if the gender of the highlighted word traduttore is changed to the feminine traduttrice, (edit correctly and fully predicted) but since word stem characters are left unchanged. Table 1 (Measured) shows metric values across the three highlight modalities (breakdowns by domain and speed shown in Tables 10 and 11). As expected, Oracle highlights obtain the best performance in terms of precision and recall, with , in particular, being significantly higher than the other two modalities across both directions.
Surprisingly, we find no significant precision and recall differences between Supervised and Unsupervised highlights despite the word-level QE training of XCOMET used in the former modality. Moreover, they support the potential of unsupervised, model internals-based techniques to complement or substitute more expensive supervised approaches. Still, likelihood ratios for all modalities and directions indicate that highlights are 2-4 times more likely to precisely and comprehensively encompass edits than non-highlighted texts. This suggests that even imperfect highlights that do not reach Oracle-level quality might effectively direct editing efforts toward potential issues. We validate these observations by fitting a zero-inflated negative binomial mixed-effects model to predict segment-level edit rates. Results confirm a significantly higher edit rate for EnglishItalian highlighted modalities and the social media domain with (features and significances shown in Appendix Table 6). We find a significant zero inflation associated with translator identity, suggesting the choice of leaving MT outputs unedited is highly subjective.
@@
Having observed a clear influence of highlights on the editing rate of translators, we question whether the relatively high and values for some highlight modalities are not artificially inflated by the eagerness of translators to intervene on highlighted spans, rather than reflecting actual highlight quality. In other words, do highlights identify actual issues, or do they condition translators to edit when they otherwise would not? To answer this, we propose to project highlights from a selected modality—in which highlights were shown during editing—onto the edits performed by the No Highlight translators on the same segments. The resulting difference between measured and projected metrics can then be taken as an estimate for the impact of highlight presentation on their resulting accuracy. To further ensure the soundness of our analysis, we use a set of projected Random highlights as a lower bound for highlight performance. To make the comparison fair, Random highlights are created by randomly highlighting words in MT outputs matching the average word-level highlight frequency across all highlighted modalities given the current domain and translation direction. Table 1 (Projected) shows results for the three highlighted modalities. Projected precision scores are influenced by edit frequency, and hence see a major decrease for EnglishItalian, where the No Highlight edit rate is much lower. However, the increase in across all EnglishItalian modalities confirms that, despite the lower edit proportion, highlighted texts remain notably more likely to be edited than non-highlighted ones. Conversely, the lower for EnglishDutch shows that edits become much less skewed towards highlighted spans in this direction when accounting for presentation bias. and , which are not influenced by different edit rates, exhibit the same trend. Finally, projected metrics remain consistently above the Random baseline, suggesting a higher-than-chance ability to identify errors even in the worst-performing settings.
@@ 4.3 Quality Assessment
We continue our assessment by inspecting the quality of MT and post-edited outputs along three dimensions. First, we use XCOMET segment-level QE ratings as an automatic approximation of quality and compare them to human-annotated quality scores collected in the last phase of our study. For efficiency, these are obtained for the 0-100 Direct Assessment scale commonly used in QE evaluation (Specia et al., 2020), but following an initial step of MQM error annotation to condition scoring on found errors, as prescribed by the ESA protocol (Kocmi et al., 2024b). Secondly, MQM error span annotations are used to analyze the distribution of error categories. Thirdly, we manually assess critical errors, which we inserted to quantify whether highlight modalities can help address unambiguous issues.
@@
In this stage, we focus particularly on edited quality improvements, i.e. the quality difference between MT and post-edited versions of the same segment. Positive scores for this metric reflect an increase in quality after the post-editing stage and are bounded by the maximal achievable quality gain given the initial MT quality. Figure 4 shows median improvement values across quality bins defined from the distribution of initial MT quality scores (shown in histograms), in which all post-edited versions of each MT output appear as separate observations. Positive median scores confirm that post-edits generally lead to quality improvements across all tested settings. However, we observe different trends across the two metrics: across both domains, XCOMET underestimates the human-assessed ESA quality improvement, especially for biomedical texts where it shows no improvements regardless of the initial MT quality. These results echo recent findings cautioning users against the poor performance of trained MT metrics for unseen domains and high-quality translations (Agrawal et al., 2024; Zouhar et al., 2024a). Focusing on the more reliable ESA scores, we observe large quality improvements from post-editing, as shown by near-maximal quality gains across most bins and highlight modalities. While No Highlight seems to underperform other modalities in the social media domain, the lack of more notable differences suggests that highlights’ quality impact might not be evident in terms of segment-level quality, motivating our next steps in the quality analysis.
We also find no clear relationship between translator speed and edited quality improvements, suggesting that higher productivity does not come at a cost for faster translators (Figure 10). This finding further confirms that neglecting errors is not the cause of the differences in editing patterns shown in previous sections.
@@
Figure 5 shows a breakdown of MQM annotations across modalities, domains and directions for three error macro-categories.777Detailed MQM error counts are shown in Table 9. At this granularity, differences across modalities become visible, with overall error counts showing a clear relation to from Table 1 (Oracle being remarkably better for EnglishItalian, with milder and more uniform trends in EnglishDutch). At least for EnglishItalian, these results confirm that an observable quality improvement from editing with highlights is present in the best-case Oracle scenario. We also note a different distribution of Accuracy and Style errors, with the former being more common in biomedical texts while the latter appearing more often for translated social media posts. We posit that differences in error types across domains might be the culprit behind the opposite productivity trends observed in Section 4.1: highlighted accuracy errors might lead to time-consuming terminology checks in biomedical texts. In contrast, style errors with simple fixes might be corrected more quickly in social media texts.
@@
We examine whether the critical errors we inserted were detected by different modalities, finding that while most modalities fare decently with more than 62% of critical errors highlighted, Unsupervised is the only setting for which all errors are correctly highlighted across both directions. Then, critical errors are manually verified in all outputs, finding that 16-20% more critical errors are edited in highlighted modalities compared to No Highlight (full results in Table 9). Hence, the presence of highlights might result in narrow but tangible quality improvements that remain undetected in coarser assessments, and finer-grained evaluations might be needed to quantify future improvements in word-level QE quality.
@@ 4.4 Usability
In post-task questionnaire answers (Table 2), most translators stated that MT outputs had average-to-high quality and that provided texts were challenging to translate. Highlights were generally found decently accurate, but they were generally not found useful to improve either productivity or quality (including Oracle ones). Interestingly, despite the convincing gains for critical errors measured in the last section, most translators stated that highlights did not influence their editing and did not help them identify errors that would have otherwise been missed. Concretely, this suggests that the potential quality improvements might not be easily perceived by translators and might have secondary importance compared to the extra cognitive load elicited by highlighted spans. When asked to comment about highlights, several translators called them “more of an eye distraction, as they often weren’t actual mistakes” and “not quite accurate enough to rely on them as a suggestion”. Some translators also stated that missed errors led them to “disregarding the highlights to focus on checking each sentence”. Despite their high quality, only one editor working with Oracle highlights found highlights helpful in “making the editing process faster and somehow easier”. Taken together, these comments convincingly point to a negative perception of the quality and usefulness of highlights, suggesting that improvement in QE accuracy may not be sufficient to improve QE usefulness in editors’ eyes.
| Question | EngIta | EngNld |

### v2

@@ 3.4 Participants
For both directions, professional translation companies Translated Srl555https://translated.com and Global Textware666https://globaltextware.nl/ recruited three translators for the Oracle post-editing stage, the core set of 12 translators working on Pre, Main and Post tasks, and six more translators for the QA stage, for a total of 21 translators per direction. All translators were freelancers with native proficiency in their target language and self-assessed proficiency of at least C1 in English. Almost all translators had more than two years of professional translation experience and regularly post-edited MT outputs (details in Table˜10).
@@ 3.5 Editing Interface
We develop a custom interface, which we name Groningen Translation Environment (GroTE, Figure˜2), to support editing over texts with word-level highlights. While the MMPE tool used by Shenoy et al. (2021) provide extensive multimodal functionalities (Herbig et al., 2020), we aim for a bare-bones setup to avoid confounders in the evaluation. GroTE is a web interface based on Gradio (Abid et al., 2019) and hosted on the Hugging Face Spaces to enable multi-user data collection online. Upon loading a document, source texts and MT outputs for all segments are presented in two columns following standard industry practices. For modalities with highlights, the interface provides an informative message and supports the removal of all highlights on a segment via a button, with highlights on words disappearing automatically upon editing, as in Shenoy et al. (2021). The interface supports real-time logging of user actions, allowing for the analysis of the editing process. In particular, we log the start/end times for each document, the accessing and exiting of segment textboxes, highlight removals, and keystrokes.
GroTE intentionally lacks standard features such as translation memories, glossaries, and spellchecking to ensure equal familiarity among translators, ultimately controlling for editor proficiency with these tools, as done in previous studies (Shenoy et al., 2021; Sarti et al., 2022). While most translators noted the lack of features in our usability assessment, the majority also found the interface easy to set up, access, and use (Table˜10).
@@ 4.1 Productivity
We obtain segment- and document-level edit times and compute editing productivity as the number of processed source characters over the sum of all document-level edit times, measured in characters per minute. To account for potential breaks taken by post-editors during editing, we filter out pauses between logged actions longer than 5 minutes. We note that this procedure does not significantly impact the overall ranking of translators, while ensuring a more robust evaluation of editing time.
@@
Figure˜3 shows translators’ productivity across stages, with every dot corresponding to the productivity of a single individual. We observe that no highlight modality leads to systematically faster editing across all speed groups and that the ordering of Pre-task speed groups is maintained in the following stages despite the different highlight modalities. These results suggest that individual variability in editing speed is more critical than highlight modality in predicting editing speed. However, faster EnglishDutch translators achieve outstanding productivity, i.e. standard deviations above the overall mean ( char/min, ➔ in Figure˜3) almost exclusively in No Highlight, and, Oracle modalities, suggesting that lower-quality highlights hinder editing speed.
We validate these observations by fitting a negative binomial mixed-effect model on segment-level editing times (model details in Table˜8). Excluding random factors such as translator and segment identity from the model produces a significant drop in explained variance, confirming the inherent variability of editing times (). Model coefficients show that MT output length and the proportion of highlighted characters are the main factors driving an increase in editing times, possibly reflecting an increase in cognitive effort to process additional information. We find highlights to have a significant impact on increasing the editing speed of EnglishItalian translators , but a minimal impact for EnglishDutch. Comparing the productivity of the same translator editing with and without highlights (Main vs Post), two-thirds of the translators editing with highlights were up to two times slower on biomedical texts. However, the same proportion of translators was up to three times faster on social media texts across both directions.
@@ 4.2 Highlights and Edits
First, we quantify how different modalities agree in terms of highlight distribution and editing. We find that highlight overlaps across modalities range between 15% and 39% when comparing highlight modalities in a pairwise fashion, with the highest overlap for EnglishItalian social media and EnglishDutch biomedical texts.777Scores are normalized to account for highlight frequencies across modalities. Agreement is shown in Table 11. Despite the relatively low highlight agreement, we find an average agreement of 73% for post-edited characters across modalities. This suggests that edits are generally uniform regardless of highlight modalities and are not necessarily restricted to highlighted spans.888Editing agreement is shown in Figure 7.
@@
Table˜4 (Base Freq.) shows raw highlight and edit frequencies across modalities. We observe different trends across the two language pairs: for EnglishItalian, post-editors working with highlights edit more than twice as much as translators with No Highlight, regardless of the highlight modality. On the contrary, for EnglishDutch they edit 33% less in the same setting. These results suggest a different attitude towards acting upon highlighted potential issues across the two translation directions, with EnglishItalian translators appearing to be conditioned to edit more when highlights are present. We introduce four metrics to quantify highlights-edits overlap:
@@
Intuitively, character-level recall should be more indicative of highlight quality compared to precision , provided that word-level highlights can be useful even when not minimal.999For example, if the fully-highlighted word traduttore is changed to its feminine version traduttrice, (edit correctly and fully predicted) but since word stem characters are left unchanged. Table˜4 (Measured) shows metric values across the three highlight modalities (breakdowns by domain and speed shown in Tables˜13 and 14). As expected, Oracle highlights obtain the best performance in terms of precision and recall, with , in particular, being significantly higher than the other two modalities across both directions.
Surprisingly, we find no significant precision and recall differences between Supervised and Unsupervised highlights, despite the word-level QE training of XCOMET used in the former modality. Moreover, they support the potential of unsupervised, model internals-based techniques to complement or substitute more expensive supervised approaches. Still, likelihood ratios for all modalities and directions indicate that highlights are 2-4 times more likely to precisely and comprehensively encompass edits than non-highlighted texts. This suggests that even imperfect highlights that do not reach Oracle-level quality might effectively direct editing efforts toward potential issues. We validate these observations by fitting a zero-inflated negative binomial mixed-effects model to predict segment-level edit rates. Results confirm a significantly higher edit rate for EnglishItalian highlighted modalities and the social media domain with (features and significances shown in Appendix Table˜9). We find a significant zero inflation associated with translator identity, suggesting the choice of leaving MT outputs unedited is highly subjective.
@@
Since in Section˜4.1 we found the proportion of highlighted characters to impact the editing rate of translators, we question whether the relatively high and values might be artificially inflated by the eagerness of translators to intervene on highlighted spans. In other words, do highlights identify actual issues, or do they condition translators to edit when they otherwise would not? To answer this, we propose to project highlights from a selected modality—in which highlights were shown during editing—onto the edits performed by the No Highlight translators on the same segments. The resulting difference between measured and projected metrics can then be taken as an estimate for the impact of highlight presentation on their resulting accuracy.
To further ensure the soundness of our analysis, we use a set of projected Random highlights as a lower bound for highlight performance. To make the comparison fair, Random highlights are created by randomly highlighting words in MT outputs matching the average word-level highlight frequency across all highlighted modalities given the current domain and translation direction. Table˜4 (Projected) shows results for the three highlighted modalities. First, all projected metrics remain consistently above the Random baseline, suggesting a higher-than-chance ability to identify errors even for worst-performing highlight modalities. Projected precision scores depend on edit frequency, and hence see a major decrease for EnglishItalian, where the No Highlight edit rate is much lower. However, the increase in across all EnglishItalian modalities confirms that, despite the lower edit proportion, highlighted texts remain notably more likely to be edited than non-highlighted ones. Conversely, the lower , and for EnglishDutch show that edits become much less skewed towards highlighted spans in this direction when accounting for presentation bias.
@@ 4.3 Quality Assessment
We continue our assessment by inspecting the quality of MT and post-edited outputs along three dimensions. First, we use XCOMET segment-level QE ratings as an automatic approximation of quality and compare them to human-annotated quality scores collected in the last phase of our study. For efficiency, these are obtained for the 0-100 Direct Assessment scale commonly used in QE evaluation (Specia et al., 2020), but following an initial step of MQM error annotation to condition scoring on found errors, as prescribed by the ESA protocol (Kocmi et al., 2024b). Then, MQM error span annotations are used to analyze the distribution of error categories. Finally, we manually assess critical errors, which were inserted to quantify highlight modalities effect on unambiguous issues.
@@
In this stage, we focus particularly on edited quality improvements, i.e. how post-editing the same MT outputs under different highlight conditions influences the resulting quality of translations. We operationalize this assessment using human ratings and automatic metrics to score MT and post-edited translations, using their difference as the effective quality gain after the post-editing stage. Scores for this metric are generally positive, i.e. human post-editing improves quality, and bounded by the maximal achievable quality gain given the initial MT quality. Figure˜4 shows median improvement values across quality bins defined from the distribution of initial MT quality scores (shown in histograms), in which all post-edited versions of each MT output appear as separate observations. Positive median scores confirm that post-edits generally lead to quality improvements across all tested settings. However, we observe different trends across the two metrics: across both domains, XCOMET greatly underestimates the human-assessed ESA quality improvement, especially for biomedical texts where it shows negligible improvement regardless of the initial MT quality. These results echo recent findings cautioning users against the poor performance of trained MT metrics for unseen domains and high-quality translations (Agrawal et al., 2024; Zouhar et al., 2024). Focusing on the more reliable ESA scores, we observe large quality improvements from post-editing, as shown by near-maximal quality gains across most bins and highlight modalities. While No Highlight seems to underperform other modalities in the social media domain, the lack of more notable differences in gains across highlight modalities suggests that highlights’ quality impact might not be evident in terms of segment-level quality, motivating our next steps in the quality analysis.
We also find no clear relationship between translator speed and edited quality improvements, suggesting that higher productivity does not come at a cost for faster translators (Figure˜10). This finding confirms that neglecting errors is not the cause of different editing patterns from previous sections.
| MT | No High. | Oracle | Unsup. | Sup. | ||
| Italian | Acc. | 26 / 31 | 12 / 17 | 6 / 12 | 30 / 22 | 22 / 24 |
| Style | 17 / 33 | 5 / 33 | 0 / 15 | 5 / 35 | 4 / 31 | |
| Ling. | 12 / 31 | 2 / 29 | 0 / 11 | 7 / 20 | 2 / 16 | |
| Tot. | 55 / 95 | 19 / 79 | 6 / 38 | 42 / 77 | 28 / 71 | |
| Dutch | Acc. | 32 / 40 | 20 / 31 | 26 / 37 | 14 / 39 | 20 / 39 |
| Style | 3 / 29 | 1 / 28 | 1 / 23 | 2 / 18 | 7 / 50 | |
| Ling. | 4 / 26 | 3 / 18 | 5 / 28 | 2 / 9 | 3 / 15 | |
| Tot. | 39 / 95 | 24 / 77 | 32 / 88 | 18 / 66 | 30 / 104 |
@@
Table˜5 shows a breakdown of MQM annotations for MT and all highlight modalities using the Accuracy, Style and Linguistic macro-categories of MQM errors.101010Full micro-category breakdown in Table 12, per-domain breakdown in Figure 9. Category descriptions in Figure 5. At this granularity, differences across modalities become visible, with overall error counts showing a clear relation to from Table˜4 (Oracle being remarkably better for EnglishItalian, with milder and more uniform trends in EnglishDutch). At least for EnglishItalian, these results confirm that an observable quality improvement from editing with highlights is present in the best-case Oracle scenario. By contrast, for EnglishDutch the Unsupervised method is found to outperform even the Oracle setting in reducing the amount of errors, while it fares relatively poorly for EnglishItalian. We also note a different distribution of Accuracy and Style errors, with the former being more common in biomedical texts while the latter appearing more often for translated social media posts (Figure˜9). We posit that differences in error types across domains might explain the opposite productivity trends observed in Section˜4.1: while highlighted accuracy errors might lead to time-consuming terminology verification in biomedical texts, style errors might be corrected more quickly and naturally in the social media domain.
@@
We examine whether the critical errors we inserted were detected by different modalities, finding that while most modalities fare decently with more than 62% of critical errors highlighted, Unsupervised is the only setting for which all errors are correctly highlighted across both directions. Then, critical errors are manually verified in all outputs, finding that 16-20% more critical errors are edited in highlighted modalities compared to No Highlight (full results in Table˜12). Hence, highlights might lead to narrow but tangible quality improvements that can go undetected in coarse quality assessments, and finer-grained evaluations might be needed to quantify future improvements in word-level QE.
@@ 4.4 Usability
In post-task questionnaire answers (Table˜6), most translators stated that MT outputs had average-to-high quality and that provided texts were challenging to translate. Highlights were generally found decently accurate, but they were generally not found useful to improve either productivity or quality (including Oracle ones). Interestingly, despite the convincing gains for critical errors measured in the last section, most translators stated that highlights did not influence their editing and did not help them identify errors that would have otherwise been missed. Concretely, this suggests that the potential quality improvements might not be easily perceived by translators and might have secondary importance compared to the extra cognitive load elicited by highlighted spans. When asked to comment about highlights, several translators called them “more of an eye distraction, as they often weren’t actual mistakes” and “not quite accurate enough to rely on them as a suggestion”. Some translators also stated that missed errors led them to “disregarding the highlights to focus on checking each sentence”. Despite their high quality, only one editor working with Oracle highlights found highlights helpful in “making the editing process faster and somehow easier”. Taken together, these comments convincingly point to a negative perception of the quality and usefulness of highlights, suggesting that improvement in QE accuracy may not be sufficient to improve QE usefulness in editors’ eyes.
| Question | Italian | Dutch |
@@ 5 Conclusion
6 Limitations
Our study presents certain limitations that warrant consideration when interpreting its findings and for guiding future research.
Firstly, while we included two domains and translation directions to improve the generalizability of our findings, our results suggest that language and domain play an important role in defining the effectiveness of word-level QE for human post-editing. While we observed mild gains from word-level QE on our tested mid-resourced translation directions (EnglishItalian and EnglishDutch), we expect limited, if any, benefit of such approaches in low-resource languages and domains for which MT systems and QE methods are likely to underperform (Sarti et al., 2022; Zouhar et al., 2024). Furthermore, the domains tested in our study (biomedical and social media posts) provided concrete challenges in the form of specialized terminology and idiomatic expressions, respectively, which are known to hinder the quality of MT outputs (Neves et al., 2024; Bawden and Sagot, 2023). While future work should ensure our findings can be extended to other domains and languages, the limited benefits brought by the tested word-level QE methods in challenging settings suggest a limited usefulness for higher-resource languages and more standard domains such as news or Wiki texts.
Secondly, we acknowledge that several design choices in our evaluation setup, rather than pertaining to the QE methods themselves, may have influenced our results. These include, for instance, the specific procedure for discretizing continuous scores from the Unsupervised method into error spans, and the method of obtaining oracle highlights via majority voting among post-editors. While we believe these choices are justified within the context of our study, their impact on the outcomes cannot be entirely discounted. Future studies might benefit from a more fine-grained assessment of how such low-level decisions influence the perceived accuracy and usability of word-level QE.
Finally, subjective factors such as the translators’ inherent propensity to edit, their prior opinions on the role of MT in post-editing, and their individual editing styles inevitably influenced both quantitative and qualitative assessments in this study. Although we attempted to mitigate these effects by ensuring a controlled evaluation setup for all professional translators and by using averaged judgments for translators working on the same highlight modality, we acknowledge that subjectivity might limit the reproducibility of our findings.
7 Broader Impact and Ethical Considerations
Our study explicitly centers the experience of professional translators, responding to recent calls for user-centered evaluation of translation technologies. By prioritizing translators’ perspectives and productivity, we aim to contribute to methods that complement rather than replace human expertise. Our findings highlight a gap between user perception and measured quality improvements, suggesting that future efforts should focus primarily on improving the usability of these methods in editing interfaces. In particular, new assistive approaches for post-editing should not only strive to increase productivity but rather reduce the cognitive burden associated with post-editing work. This insight is crucial for designing more user-centered quality estimation tools that genuinely support human work. Ultimately, our results suggest that subjective norms across different domains and cultures play an important role in determining the effectiveness of proposed methodologies, underscoring the importance of accounting for human factors when designing such evaluations. All participants in this study were professional translators who provided informed consent. The research protocol ensured anonymity and voluntary participation, with translators recruited and remunerated through professional translation providers. The released materials further promote transparency, enabling other researchers to reproduce and build upon our findings.